East Commercial Bank offers to lend you $40,000 at a nominal rate of 7.5%, compounded monthly. The loan (principal plus interest) must be repaid at the end of the year. West Commercial Bank also offers to lend you the $40,000, but it will charge an annual rate of 8.0%, with no interest due until the end of the year. How much higher or lower is the effective annual rate charged by West Commercial Bank versus the rate charged by East Commercial Bank?
Answer #1

EAR of East Commercial Bank =(1+APR/12)^12-1 =(1+7.5%/12)^12-1 =7.76%
EAR of West Commercial Bank =8%
EAR of West commercial bank is higher by 8%-7.76% =0.24%